23. Draw the new product Lindlars Catalyst 24. draw the products. If no reaction, put NR NaNIL ] ܢ ܕ.O ܢܗܝ . ܝܟܝ ܝܨ ܓ ܂
Answer #1

HE Hge c4g - CH, a) Hac-c=c- CH - C +, - - 2-pentyne Lindlars Cataly 87 cis -2- pentene Lindlars catalyst alkynes undergo Re-cZc-t Nartiya r = NO & N43 weak acid ich Na CEC CEC-c-H H20 € OH C-CAR (product) sc=c-H Naultz -c c -A Naoh . I No Reaction